# Improved implicit diffusion model with knowledge distillation to estimate the spatial distribution density of carbon stock in remote sensing imagery

### v1

@@ Methods
The Knowledge Distillation (KD) module consists of four parts, and a detailed description is provided in Appendix Section 2.
@@
Source model. VGG-19 was selected as an illustrative model for our approach. In this context, the input layer of VGG-19 served as the source model, denoted as $ENC$, from which the knowledge of the features was extracted. Subsequently, this knowledge was transferred to a smaller target model, referred to as $enc$. The architecture of the $enc$ model mirrored that of the source model ($ENC$), albeit with a reduced channel length at each layer.
@@
Blockwise PCA-based KD. To facilitate feature transformation within the distillation model, it was necessary to incorporate a paired decoder denoted as $dec$. This decoder worked in tandem with the encoder $enc$ to extract the input information effectively. The distillation method used in this context was Principal Components Analysis (PCA). This module is shown in Fig. 2
Figure 2: PCA-based knowledge distillation structure for VGG. Notes: PCA-based knowledge distillation consists of two steps: (a) global eigenbasis (${{\mathbf{W}}_{\mathbf{N}}},\mathbf{N}=1,2,3,...,16$) derivation, (b) blockwise PCA knowledge distillation, (c) the basic autoencoder framework of VGG-19, and (d) the $enc-dec$ autoencoder framework from our method. $relu16$ = $relu16^{\prime}$, and $relu16{{}_{e}}$ = $relu16{{}_{d}}$.
Denoising model
The conditional network used a Convolutional Neural Network (CNN) to extract conditional features, with the KD-VGG module establishing the initial features ${{\mathbf{f}}^{(0)}}$. By optimizing KD-VGG, we increased the network depth for deeper feature extraction and significantly reduced the number of parameters, enhancing performance without additional computational overhead.
Subsequently, we concatenated ${{\mathbf{f}}^{(0)}}$ with ${{\mathbf{y}}_{\mathbf{t}}}$ and fed them into the UNet for encoding. Simultaneously, they were downsampled to generate ${{\mathbf{f}}^{(i)}}$, as illustrated in Eq. 2.
@@
Where, the serial number $i$ represented the output of different layers, and $t$ was the time. In contrast to GAN-based methodologies, which depend on the incorporation of additional prior knowledge, this conditional network only offers encoded features and transmits them to the UNet, all without the need for additional prior information to model potential representations.
This study constituted an estimation task, in which the dimensions of both the input and output images remained unchanged. In particular, modifications were made to the modulation of the scaling factor ($s$) modulation within the original model. By setting this factor to 1, the zoom control for that specific component was eliminated. Building on the foundation of the original model [35], we introduced attention mechanisms to facilitate the fusion of features derived from ${{\mathbf{f}}^{(i)}}$ and ${{\mathbf{u}}^{(i)}}$, followed by their input into the implicit representation module.
Implicit representation
Presently, prevalent algorithms often rely on intricate cascading or two-stage training procedures. However, our research discovered that the use of implicit neural representations for image representation learning notably enhanced estimation accuracy. To implement this, we incorporated several coordinate-based Multi-layer Perceptrons (MLPs) into the up-sampling stage of the UNet architecture, thereby enabling the parameterization of implicit neural representations. We employed a two-layer MLP denoted as $D$ to perform up-sampling.
@@
Improved implicit diffusion model (IIDM)
The Implicit Diffusion Model (IDM) combined diffusion models and implicit neural representations for image-to-image transformation. The Improved Implicit Diffusion Model (IIDM) workflow is shown in Fig. 3. We used a Recurrent Neural Network (RNN) to create a time-dependent structure with shared parameters, including a denoising and implicit representation model, employing variational inference. Unlike Variational Autoencoders (VAE), the implicit variables in diffusion models have the same dimensionality as the original data, with a predefined and fixed diffusion process.
The Diffusion principle’s derivation of forward and reverse processes is shown in Appendix Section 3. The diffusion model included two processes: the diffusion process (that is, the forward process) and the reverse process. Both the forward and reverse processes were a parameterized Markov chain, in which the reverse process can be used to generate data.
The added KD-VGG module can extract the shallow features of $\mathbf{x}$ and $\mathbf{y_{t}}$, which are then passed into KD-UNet for encoding and decoding to reconstruct the spatial distribution density image. The concatenation of $\mathbf{f^{0}}$ and $\mathbf{y_{t}}$ enables real-time updating of the model, i.e., updating with the change of $t$. Cross-attention effectively handles the relationship between the two features, facilitating their fusion. The multi-layer perceptron (MLP) merges the condition and features, increasing the influence of the condition on the parameters and speeding up the fitting process.
@@
The implicit diffusion model aimed to infer the target image $\mathbf{y_{0}}$ using a series of denoising steps. To achieve this, the denoising model ${{\epsilon}_{\theta}}$ was optimized to restore the target image $\mathbf{y_{0}}$ from a noisy target image ${{\mathbf{\tilde{y}}}_{\mathbf{t}}}=\sqrt{{{\gamma}_{t}}}{{\mathbf{y}}_{\mathbf{0}}}+\sqrt{1-{{\gamma}_{t}}}{\epsilon}$. To achieve this, the denoising model ${{\epsilon}_{\theta}}$ was optimized to restore the target image $\mathbf{y_{0}}$ from a noisy target image ${{\mathbf{\tilde{y}}}_{\mathbf{t}}}=\sqrt{{{\gamma}_{t}}}{{\mathbf{y}}_{\mathbf{0}}}+\sqrt{1-{{\gamma}_{t}}}{\epsilon}$. Ultimately, the denoising network was optimized to achieve the goal while maintaining the accuracy of the predicted noise ${\epsilon}$, as shown in Eq. 4.
@@
Mean Absolute Error (MAE) measures the average absolute error between predicted and actual values. Mean Squared Error (MSE) is the average squared error, and Root Mean Squared Error (RMSE) is the square root of the MSE. Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Metric (SSIM) are full-reference image quality indices, with PSNR assessing overall image quality and SSIM evaluating similarity in brightness, contrast, and structure.
Results
Inference time
@@ Results
We used $256\times 256$ images, and the test results included seven models (as shown in Appendix Table A3): VGG, KD-VGG, Stable Diffusion (SD), DDPM, DDIM, Ours-VGG, and Ours-KD-VGG. DDPM, DDIM, and SD were diffusion models with more than 450M parameters. Adding modules for improved accuracy could lead to high system overhead and reduced feasibility, prompting a discussion on the justification for such additions.
The VGG-19 model, originally with 76.39M parameters, was reduced to 0.99M parameters after dimensionality reduction, achieving approximately 77 times compression. The KD-VGG-19 model retained most of VGG-19’s feature extraction capabilities while significantly reducing its size, making it a lightweight and efficient feature extraction module for downstream tasks. Similar compression methods were applied to VGG-16 and VGG-11, achieving compression ratios of 89.49 and 184.80 times, respectively. Among the three, VGG-11 had the highest compression efficiency, followed by VGG-16, while VGG-19, though least compressed, performed best in estimation.
@@
The VGG module was incorporated into the feature extraction section of IDM. KD-VGG maintained the model’s robust feature extraction capabilities and improved estimation accuracy while reducing inference time with scaled-down parameters. The inference times for KD-VGG-11, KD-VGG-16, and KD-VGG-19 were 97.22%, 86.22%, and 89.97% of their pre-KD times, respectively, while the inference time of KD-UNet was 61.92% of the original UNet. UNet compression had a more significant impact on inference time compared to VGG. The IDM parameters were only about 9% of the SD’s, with 11.28M more than DDPM. Our VGG model required 72.80$\sim$73.60M fewer parameters than IDM, accounting for just $\sim$7.5of SD. DDPM had the fastest inference time among the diffusion models, and our VGG model was even faster than IDM by 0.17$\sim$0.18 seconds. The reduction in model parameters enables easy deployment in common devices and the reduced inference time facilitates the application of large models to quantitative remote sensing with substantial data volumes.
Model compression
KD-VGG.
During the distillation process, principal components with a cumulative contribution rate higher than 85% were selected to maximize effective information extraction unsupervisedly. This enabled channel dimensionality reduction, decreasing model parameters. VGG initially presented an increasing channel length, which was reduced after dimensionality reduction. This indicated that deeper layers had more invalid channels, and reducing these channels effectively decreased parameters with minimal impact on accuracy. This is likely because most effective features are extracted in the middle layers, while deeper layers do not require such complexity. VGG-11, VGG-16, and VGG-19 all showed improved performance with deeper layers. For example, VGG-19’s distillation results are shown in Table A2, and the dimensionality reduction results for 16 features are depicted in Figure 4.
@@
KD-UNet.
PCA performed dimensionality reduction based on feature transformations and spatial mapping, while PCA-based distillation was interpretable and generalizable and can be used for model compression of other network structures. IDM contained the UNet structure, and we applied this to the UNet with a view to maximize the compression of diffusion models. The distillation results are shown in Appendix Table A2, and the channel downscaling results for the 10 features are shown in Appendix Fig. A1. UNet was a $enc-dec$ structure and requires a skip connection, which required the same number of channels in the front and back structure. Moreover, there existed an upsample between each 2 layers, so it was necessary to present a double number of channels. We added Selected column in Appendix Table A2 to describe the final number of selected channels. The selection was based on the minimum channel length that meets the requirements, and the final selection was (44, 44, 88, 88, 176, 176, 352, 352, 704, 704).
Ablation study
Five modules were selected for the ablation experiment: Mask, VGG, KD-VGG, KD-UNet and Attention + MLP, detailed in Table 1. Mask filtered vegetation and non-vegetation areas effectively, serving as the initial data preprocessing stage. VGG and KD-VGG modules were employed for the initial extraction of features, with a comparable performance observed between KD-VGG and VGG models. However, VGG slightly outperformed KD-VGG, highlighting the effectiveness of knowledge distillation in extracting critical model parameters while indicating areas for further enhancement. Introducing attention and MLP modules resulted in improved model performance, albeit with increased computational influence time. The results indicated a significant accuracy enhancement after the integration of the modules, confirming the optimization achieved through the joint use of the KD-VGG and Attention + MLP modules. Thus, the strategy of prioritizing accuracy over computational time proved beneficial.
Table 1: Compare of modules, including Mask, VGG, KD-VGG, and Attention + MLP. Notes: We choose VGG-19 and KD-VGG-19. Bold is the best, and underline is the second. KD-V = KD-VGG, KD-U = KD-UNet, and A+MLP = Attention + MLP.
@@
1 $\times$ $\checkmark$ $\times$ $\times$ $\times$ 0.0965 0.1670 0.5986 19.0697
2 $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ 0.0924 0.1697 0.6098 18.9279
3 $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ 0.0921 0.1680 0.6357 19.0166
4 $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ 0.0924 0.1692 0.6073 18.9546
5 $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 0.0925 0.1684 0.6192 18.9969
6 $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ 0.0959 0.1670 0.5879 19.0662
7 $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ 0.0742 0.1363 0.7283 20.8318
8 $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ 0.0692 0.1229 0.7196 21.7321
9 $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ 0.0687 0.1211 0.7289 21.8581
10 $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ 0.0691 0.1288 0.7388 21.4396
11 $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 0.0688 0.1217 0.7186 21.8167
12 $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ 0.0702 0.1306 0.7207 21.2021
Comparison of different models
We selected seven comparison models, including commonly used regression models in remote sensing estimation (OLS, RF, and SVR), and deep learning generative models (VAE, GAN, and IIDM proposed in this paper). The results are presented in Table 2. OLS was the simplest and least computationally expensive algorithm among all models. RF and SVR were the most widely used machine learning algorithms for estimation. However, the accuracy of these three algorithms was similar and generally poor in all models. The accuracy of deep learning algorithms was significantly better, with generative models outperforming VAE and diffusion models outperforming GAN. Among all models, IDM-VGG achieved the best performance, followed closely by IDM-KD-VGG (IIDM), with only a slight difference in accuracy between the two. Taking into account both the parameters and the inference time, IIDM proved to be more versatile.

### v2

@@ 4.3 Knowledge distillation (KD) module
The Knowledge Distillation (KD) module consists of four parts, and a detailed description is provided in Appendix Section B.
@@
Source model. VGG-19 was selected as an illustrative model for our approach. In this context, the input layer of VGG-19 served as the source model, denoted as $ENC$, from which the knowledge of the features was extracted. Subsequently, this knowledge was transferred to a smaller target model, referred to as $enc$. The architecture of the $enc$ model mirrored that of the source model ($ENC$), albeit with a reduced channel length at each layer. This reduction helps maintain computational efficiency while preserving critical feature extraction capabilities.
@@
Blockwise PCA-based KD. To facilitate feature transformation within the distillation model, it was necessary to incorporate a paired decoder denoted as $dec$. This decoder worked in tandem with the encoder $enc$ to extract the input information effectively. The distillation method used in this context was Principal Components Analysis (PCA). PCA helps identify principal components that retain the most variance, effectively reducing dimensionality while preserving essential structural information. This module is shown in Figure 3.
Reducing channel lengths. In accordance with the empirical principles of PCA dimensionality reduction, it was imperative to preserve the most vital information encapsulated in the channel length $L_{N}^{e}$ of the target model $enc$. Specifically, the target layer $reluN{{}_{e}}$ of $L_{N}^{e}$ should retain variance information exceeding 85% of that found in the source layer $reluN$. This ensures that the reduced model maintains high representational capacity while significantly lowering computational complexity.
Figure 3: PCA-based knowledge distillation framework for VGG. Notes: The PCA-based knowledge distillation process consists of four key steps: (a) global eigenbasis derivation (${{\mathbf{W}}_{\mathbf{N}}},\mathbf{N}=1,2,3,...,16$), (b) blockwise PCA knowledge distillation for feature compression and transfer, (c) the standard autoencoder framework of VGG-19, and (d) the proposed $enc-dec$ autoencoder framework, which enhances efficiency while preserving key structural information. In our method, $relu16$ corresponds to $relu16^{\prime}$, while the distilled feature representation is denoted as $relu16{{}_{e}}=relu16{{}_{d}}$.
4.4 Denoising model
The conditional network utilized a Convolutional Neural Network (CNN) to extract conditional features, with the KD-VGG module establishing the initial features ${{\mathbf{f}}^{(0)}}$. By leveraging KD-VGG, the network depth was increased to enhance feature extraction capability, while the parameter count was significantly reduced, maintaining computational efficiency without compromising performance.
Subsequently, ${{\mathbf{f}}^{(0)}}$ was concatenated with ${{\mathbf{y}}_{\mathbf{t}}}$ and fed into the UNet for encoding. Simultaneously, they were downsampled to generate ${{\mathbf{f}}^{(i)}}$, as illustrated in Eq. 2.
@@ 4.4 Denoising model
Here, the index $i$ represents the different network layers, while $t$ denotes the time step within the diffusion process. Unlike GAN-based approaches, which require additional prior knowledge, this conditional network provides encoded features directly to the UNet, eliminating the need for explicit prior assumptions in potential representation modeling.
This study focused on an estimation task, where both the input and output image dimensions remained unchanged. To ensure structural consistency, modifications were applied to the scaling factor ($s$) within the original model. By setting this factor to 1, scaling adjustments were effectively neutralized, preserving spatial fidelity. Building on the foundation of the original model Gao et al. (2023b), we introduced attention mechanisms to facilitate the fusion of features extracted from ${{\mathbf{f}}^{(i)}}$ and ${{\mathbf{u}}^{(i)}}$, enhancing information aggregation before passing it into the implicit representation module.
4.5 Implicit representation
Traditional image estimation algorithms often rely on complex multi-stage training or cascading frameworks, which can lead to increased computational overhead and inefficiencies. However, our research demonstrates that the use of implicit neural representations for image representation learning can significantly enhance both estimation accuracy and model efficiency.
To implement this, we incorporated several coordinate-based Multi-layer Perceptrons (MLPs) into the up-sampling stage of the UNet architecture, allowing the network to model fine-grained spatial relationships with fewer explicit constraints. This approach enables the parameterization of implicit neural representations, effectively capturing structural and contextual dependencies in the data. We employed a two-layer MLP denoted as $D$ to perform up-sampling, as described in Eq. 3.
@@ 4.5 Implicit representation
4.6 Improved implicit diffusion model (IIDM)
The Implicit Diffusion Model (IDM) integrates diffusion models and implicit neural representations for image-to-image transformation. The Improved Implicit Diffusion Model (IIDM) workflow is shown in Figure 4. To enhance temporal consistency and computational efficiency, we used a Recurrent Neural Network (RNN) to create a time-dependent structure with shared parameters, including a denoising and implicit representation model, which employs variational inference for robust feature learning. Unlike Variational Autoencoders (VAE), the implicit variables in diffusion models have the same dimensionality as the original data, allowing for a predefined and fixed diffusion process, ensuring stability during training.
The theoretical foundation of the diffusion principle’s derivation of forward and reverse processes is shown in Appendix Section C. The diffusion model included two processes: the diffusion process (that is, the forward process) and the reverse process. Both the forward and reverse processes were modeled as a parameterized Markov chain, where the reverse process serves as a generative mechanism to reconstruct high-fidelity images.
The integration of the KD-VGG module allows for the efficient extraction of shallow image features from $\mathbf{x}$ and $\mathbf{y_{t}}$, which are then passed into KD-UNet for encoding and decoding to reconstruct the spatial distribution density image. The concatenation of $\mathbf{f^{0}}$ and $\mathbf{y_{t}}$ enables real-time updating of the model, i.e., updating with the change of $t$. Additionally, the cross-attention mechanism enhances the relationship between encoded features, facilitating improved feature fusion. The multi-layer perceptron (MLP) merges the condition and features, strengthening the impact of learned representations on the final output while accelerating the convergence of the training process.
@@ 4.6 Improved implicit diffusion model (IIDM)
The implicit diffusion model aimed to infer the target image $\mathbf{y_{0}}$ using a series of denoising steps. To achieve this, the denoising model ${{\epsilon}{\theta}}$ was optimized to restore the target image $\mathbf{y_{0}}$ from a noisy target image ${{\mathbf{\tilde{y}}}{\mathbf{t}}}=\sqrt{{{\gamma}{t}}}{{\mathbf{y}}{\mathbf{0}}}+\sqrt{1-{{\gamma}_{t}}}{\epsilon}$. This iterative process refines the estimation at each step, progressively reducing noise while preserving structural fidelity. Ultimately, the denoising network was optimized to achieve the goal while maintaining the accuracy of the predicted noise ${\epsilon}$, as shown in Eq. 4.
@@
Mean Absolute Error (MAE) measures the average absolute error between predicted and actual values. Mean Squared Error (MSE) is the average squared error, and Root Mean Squared Error (RMSE) is the square root of the MSE. Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Metric (SSIM) are full-reference image quality indices, with PSNR assessing overall image quality and SSIM evaluating similarity in brightness, contrast, and structure. These metrics collectively provide a comprehensive evaluation of model performance, balancing precision, noise resilience, and perceptual similarity.
5 Results
5.1 Inference time
@@
We used $256\times 256$ images, and the test results included seven models (as shown in Appendix Table A3): VGG, KD-VGG, Stable Diffusion (SD), DDPM, DDIM, Ours-VGG, and Ours-KD-VGG. DDPM, DDIM, and SD were diffusion models with more than 450M parameters. Adding modules for improved accuracy could lead to high system overhead and reduced feasibility. This section evaluates the balance between model complexity and computational efficiency, highlighting IIDM’s advantages in reducing inference time while maintaining high accuracy.
Using knowledge distillation, the VGG-19 model’s parameters were reduced by over 98%, significantly enhancing computational efficiency while preserving key feature extraction capabilities. Through KD-based compression, IIDM achieves a 30% reduction in inference time compared to conventional diffusion models, making it a viable approach for real-time remote sensing applications.
@@
The VGG module was incorporated into the feature extraction section of IDM. KD-VGG maintained the model’s robust feature extraction capabilities and improved estimation accuracy while reducing inference time with scaled-down parameters. The inference times for KD-VGG-11, KD-VGG-16, and KD-VGG-19 were 97.22%, 86.22%, and 89.97% of their pre-KD times, respectively, while the inference time of KD-UNet was 61.92% of the original UNet. UNet compression had a more significant impact on inference time than VGG, reducing computational cost by approximately 38%.
Unlike conventional diffusion models, IIDM integrates knowledge distillation and structured pruning to achieve a 50% reduction in computational cost, while maintaining accuracy within a 1% margin of high-parameter models. The IDM parameters were only about 9% of the SD’s, with 11.28M more than DDPM. Our VGG model required 72.80$\sim$73.60M fewer parameters than IDM, accounting for just $\sim$7.5of SD. Among the diffusion models, DDPM exhibited the fastest inference time. However, IIDM surpasses IDM in inference time by 0.17s–0.18s per image while maintaining comparable spatial representation.
By reducing model size and inference time, IIDM ensures efficient deployment on standard hardware, making it scalable for large-scale remote sensing applications.
5.2 Model compression
5.2.1 KD-VGG.
During the distillation process, principal components with a cumulative contribution rate higher than 85% were selected to maximize effective information extraction. This enabled channel dimensionality reduction, significantly decreasing model parameters while retaining crucial feature representations. Through KD-VGG compression, deeper layers were optimized by removing redundant channels, leading to a more efficient feature extraction process while preserving key spatial representations. This optimization enables IIDM to achieve a 50% reduction in model parameters while maintaining robust feature extraction capabilities.
VGG-11, VGG-16, and VGG-19 all demonstrated improved performance with deeper layers. However, through knowledge distillation, VGG-19 retained the most effective feature extraction capabilities while reducing model size. The distillation results for VGG-19 are shown in Table A2, and the dimensionality reduction for 16 features is depicted in Figure 5.
@@ 5.2 Model compression
5.2.2 KD-UNet.
To further optimize inference time, IIDM applies KD-based compression to UNet, targeting redundant channels while ensuring the integrity of skip connections. PCA performed dimensionality reduction based on feature transformations and spatial mapping, while PCA-based distillation provided interpretability and generalizability, making it applicable to model compression across different architectures.
Since UNet requires consistent channel dimensions across encoder-decoder paths, channel selection was carefully designed to preserve essential high-level features while reducing redundant layers. Maintaining structural integrity in the encoder-decoder path is crucial for diffusion models, and IIDM’s structured pruning ensures efficient information flow while achieving a 38% reduction in computational cost.
The distillation results are shown in Appendix Table A2, and the channel downscaling results for the 10 features are shown in Appendix Figure A1. The final selection of channels was based on the minimum channel length that meets structural requirements, with the final selection being (44, 44, 88, 88, 176, 176, 352, 352, 704, 704).
Overall, IIDM’s integration of knowledge distillation allows for up to a 50% reduction in parameters while maintaining state-of-the-art accuracy, making it an ideal choice for large-scale remote sensing tasks.
5.3 Ablation study
Five modules were selected for the ablation experiment: Mask, VGG, KD-VGG, KD-UNet, and Attention + MLP, detailed in Table 1. The Mask module effectively filtered vegetation and non-vegetation areas, serving as an essential preprocessing step. VGG and KD-VGG modules were employed for initial feature extraction, with KD-VGG achieving comparable performance to VGG while reducing parameters by over 98%, ensuring lightweight feature extraction. The ablation results indicate that while VGG retains slightly better feature extraction capabilities, KD-VGG achieves a 15% reduction in inference time with a negligible loss in accuracy. This highlights the effectiveness of knowledge distillation in extracting critical model parameters while maintaining computational efficiency.
Beyond feature extraction optimization, we further assess the impact of integrating attention-based mechanisms to enhance model performance. The introduction of Attention + MLP modules improved carbon stock estimation by effectively capturing global-local feature interactions. Integrating Attention + MLP improves SSIM by 5.6%, demonstrating its effectiveness in preserving structural details in carbon stock estimation. However, it also led to a slight increase in inference time, though the overall performance gain outweighed the computational cost.
The combination of KD-VGG and Attention+MLP improves overall model accuracy by 3.2%, demonstrating the effectiveness of integrating multiple feature enhancement techniques. These results confirm that a carefully balanced trade-off between computational efficiency and model accuracy—achieved through KD-VGG and Attention+MLP—optimizes IIDM for large-scale carbon stock estimation.
Table 1: Comparison of modules, including Mask, VGG, KD-VGG, and Attention + MLP. Notes: We choose VGG-19 and KD-VGG-19. Bold is the best, and underline is the second.
@@ 5.3 Ablation study
$\checkmark$ 0.0965 0.1670 0.5986 19.0697
$\checkmark$ $\checkmark$ 0.0924 0.1697 0.6098 18.9279
$\checkmark$ $\checkmark$ $\checkmark$ 0.0921 0.1680 0.6357 19.0166
$\checkmark$ $\checkmark$ 0.0924 0.1692 0.6073 18.9546
$\checkmark$ $\checkmark$ $\checkmark$ 0.0925 0.1684 0.6192 18.9969
$\checkmark$ $\checkmark$ 0.0959 0.1670 0.5879 19.0662
$\checkmark$ $\checkmark$ 0.0742 0.1363 0.7283 20.8318
$\checkmark$ $\checkmark$ $\checkmark$ 0.0692 0.1229 0.7196 21.7321
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.0687 0.1211 0.7289 21.8581
$\checkmark$ $\checkmark$ $\checkmark$ 0.0691 0.1288 0.7388 21.4396
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.0688 0.1217 0.7186 21.8167
$\checkmark$ $\checkmark$ $\checkmark$ 0.0702 0.1306 0.7207 21.2021
5.4 Comparison of different models
We selected seven comparison models, including commonly used regression models in remote sensing estimation (OLS, RF, and SVR), and deep learning generative models (VAE, GAN, and IIDM proposed in this paper). The results are presented in Table 2. OLS was the simplest and least computationally expensive algorithm among all models. RF and SVR were the most widely used machine learning algorithms for estimation. However, the accuracy of these three algorithms was similar and generally poor in all models. IIDM achieves a 41.69%–42.33% reduction in RMSE compared to traditional regression models, demonstrating its superior generalization ability. The accuracy of deep learning algorithms was significantly better, with generative models outperforming VAE and diffusion models outperforming GAN. Among all models, IDM-VGG achieved the best performance, followed closely by IDM-KD-VGG (IIDM), with only a slight difference in accuracy between the two. Taking into account both the parameters and the inference time, IIDM proved to be more versatile.